OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given points m= 7/8, (3,-4)

OpenStudy (anonymous):

use y-y1 = m(x-x1) y - -4 = 7/8 (x - 3) y + 4 = 7/8 (x - 3) 8y + 32 = 7x - 21 8y - 7x = - 53

OpenStudy (anonymous):

y=(7/8)x + b ----> -4=(7/8)*3 +b ----> b = -53/8 ---> y =(7/8)x -53/8